Green Ltd plans to open a new factory to produce smartphones. Forecast net revenue from the new factory is $75,000 per year for ten years. The smartphones will require a new factory that will cost $270,000. You can depreciate the full cost down to zero over the ten year life of the smartphone project. At the end of ten years you expect to be able to sell the equipment in the factory to a scrap dealer for $45,000. Selling the new smartphones will require additional working capital of $10,000 starting immediately. You expect to recover the working capital investment at the end of the ten year project. You have already spent $15,000 in research and development costs to source the most environmental-friendly smarphones. Assume the tax rate is 25% and your required return is 15% APR (compounded annually). Calculate the NPV, Average Accounting Return, and Payback Period for Green’s new factory.
Answer #1
year 0 1 2 3 4 5 6 7 8 9 10
initial investment -270000
working capital -10000
net revenue 75000 75000 75000 75000 75000 75000 75000 75000 75000 75000
depreciation -27000 -27000 -27000 -27000 -27000 -27000 -27000 -27000 -27000 -27000
operating profit 48000 48000 48000 48000 48000 48000 48000 48000 48000 48000
tax-25% 12000 12000 12000 12000 12000 12000 12000 12000 12000 12000
net profit 36000 36000 36000 36000 36000 36000 36000 36000 36000 36000
depreciation 27000 27000 27000 27000 27000 27000 27000 27000 27000 27000
recovery of working capital 10000
after tax sale value=45000*(1-.25) 33750
net operating cash flow -280000 63000 63000 63000 63000 63000 63000 63000 63000 63000 106750
present value factor at 15% = 1/(1+r)^n r = 15% 1 0.869565 0.756144 0.657516 0.571753 0.497177 0.432328 0.375937 0.326902 0.284262 0.247185
present value of cash flow = net operating cash flow*present value factor -280000 54782.61 47637.05 41423.52 36020.45 31322.13 27236.64 23684.03 20594.81 17908.53 26386.97
net present value = sum of present value of cash flow 46996.75
payback period in years = initial investment/annual cash flow 280000/63000 4.44
Average rate of return = average after tax profit/average investment 63000/145000 43.45%
average net profit 63000
average investment = (initial investment/2)+working capital 145000